# On the Provable Suboptimality of Momentum SGD in Nonstationary Stochastic Optimization

### v2

@@ 3.1 Bounds in expectation
$\mathcal{E}_{\beta}(\gamma)=\frac{192(2+\beta)^{2}}{\mu^{2}\gamma^{2}}\,\Delta^{2}+\frac{96}{\mu(1-\beta)}\,\sigma^{2}\gamma,\;\;\;\gamma_{\beta}^{\star}\in\arg\min_{\gamma\in(0,\ \mu(1-\beta)^{2}/(4L^{2})]}\mathcal{E}_{\beta}(\gamma),\;\;\;\mathcal{E}_{\beta}:=\mathcal{E}_{\beta}(\gamma_{\beta}^{\star}).$ (3.7)
@@
    Run SGD with momentum with constant stepsize $\gamma_{k}$ for $T_{k}$ iterations in epoch $k$, restarting the momentum buffer at the start of each epoch (equivalently, set $v=0$ or $\bm{\theta}_{t_{k}-1}=\bm{\theta}_{t_{k}}$ at epoch boundaries). Let $T:=\sum_{k=0}^{K-1}T_{k}$ be the total horizon. Then the final iterate satisfies
@@
3.2 High probability bounds
To obtain high-probability guarantees on the tracking error $\|\bm{\theta}_{t}-\bm{\theta}_{t}^{\star}\|$, we will need to make a standard light-tail assumptions on the gradient noise [HLP+19, LAN11, CDH21]:
There exists a constant $\sigma>0$ such that for all $t\geq 0$, $\big\|\bm{\xi}_{t+1}(\bm{\theta}_{t})\mid\mathcal{F}_{t}\big\|_{\Psi_{2}}\leq\sigma\;\;\text{a.s.}$
It should be noted that prior high-probability tracking analyses often iterate an MGF recursion for $\|\bm{\theta}_{t}-\bm{\theta}_{t}^{\star}\|^{2}$, which necessitates light-tail assumptions on the drift $\Delta_{t}:=\|\bm{\theta}_{t+1}^{\star}-\bm{\theta}_{t}^{\star}\|$ (e.g., conditional sub-exponentiality of $\Delta_{t}^{2}$). Our proof instead uses a optional stopping-time argument for weighted martingale difference sums. Proceeding in this way will also highlight some insights that are missing from previous bounds for SGD in nonstationary settings. We defer the proofs for this section to Appendix˜B.
@@ 3.2 High probability bounds
Theorem 3.5 (High probability tracking error bound for SGD).
Under ˜3.2, for all $t\in[T]$, $\gamma\leq\min\left\{\mu/L^{2},1/L\right\}$, and $\delta\in(0,1)$, the following tracking error bound holds for SGD with probability atleast $1-\delta$,
$\|\bm{\theta}_{T}-\bm{\theta}_{T}^{\star}\|^{2}\lesssim\left(1-\frac{\gamma\mu}{2}\right)^{T}\|\bm{\theta}_{0}-\bm{\theta}_{0}^{\star}\|^{2}+\frac{\mathfrak{D}_{T}}{\gamma\mu}+\Big(d\sigma^{2}\gamma^{2}+\frac{d^{2}\sigma^{4}\gamma^{3}}{\mu}\Big)\log\frac{2T}{\delta}+\Big(\frac{\sigma^{2}\gamma}{\mu}+\gamma^{2}\sigma^{2}\mathfrak{D}_{t}^{(2)}\Big)\log\frac{2T}{\delta}$
Prior high-probability guarantees for SGD in drifting environments typically control nonstationarity via a uniform drift bound, e.g., assuming light-tailed increments and summarizing drift by a single parameter $\Delta$ [CDH21]. Consequently, intermittent or localized nonstationarity (e.g., bursty regime shifts) is indistinguishable from persistent drift of magnitude $\Delta$ at all times, even though the dynamics exponentially forget older perturbations. In contrast, our bound is drift-adaptive and time-resolved and captures that (i) only recent nonstationarity drift affects the guarantee and (ii) drift amplifies stochastic fluctuations by increasing the tracking mismatch. This structure directly motivates restart/windowing/forgetting rules, which truncate or downweight the history and thereby reduce both drift accumulation and variance inflation after regime changes. As a consequence of Theorem˜3.5, we can get guarantees similar to Theorem˜3.2 by replicating the same argument. We exclude these results for SGD and SGDM for brevity.
@@
Theorem 3.6 (High probability tracking error bound for SGDM).

### v3

@@ C.2 Proof for SGDM tracking error bound
$\mathcal{E}_{\beta}(\gamma)=\frac{192(2+\beta)^{2}}{\mu^{2}\gamma^{2}}\,\Delta^{2}+\frac{96}{\mu(1-\beta)}\,\sigma^{2}\gamma,\;\;\;\gamma_{\beta}^{\star}\in\arg\min_{\gamma\in(0,\ \mu(1-\beta)^{2}/(4L^{2})]}\mathcal{E}_{\beta}(\gamma),\;\;\;\mathcal{E}_{\beta}:=\mathcal{E}_{\beta}(\gamma_{\beta}^{\star}).$ (95)
@@
    Run (SGDM) with constant stepsize $\gamma_{k}$ for $T_{k}$ iterations in epoch $k$, restarting the momentum buffer at the start of each epoch (equivalently, set $v=0$ or $\theta_{t_{k}-1}=\theta_{t_{k}}$ at epoch boundaries). Let $T:=\sum_{k=0}^{K-1}T_{k}$ be the total horizon. Then the final iterate satisfies
@@
Proof of Corollary˜C.6.
Fix any stepsize $\gamma\in(0,\gamma_{\max}]$ and define
$\rho(\gamma):=1-\frac{\gamma\mu}{2(1-\beta)}\in(0,1).$
By Corollary˜C.5 with this choice of $\rho$ and the inequality $\rho^{2(t+1)}\leq\exp(-\gamma\mu(t+1)/(1-\beta))$, we have for all $t\geq 0$:
$\mathbb{E}\|\bm{\theta}_{t+1}-\bm{\theta}_{t+1}^{\star}\|^{2}\;\leq\;\underbrace{\frac{48}{(1-\beta)^{2}}}_{\stackrel{{\scriptstyle\Delta}}{{=}}\;\;C_{\beta}}\exp\!\Big(-\frac{\mu\gamma}{1-\beta}(t+1)\Big)\,\|\bm{\theta}_{0}-\bm{\theta}_{0}^{\star}\|^{2}\;+\;\mathcal{E}_{\beta}(\gamma).$ (96)
Since the assumptions are uniform in time (same $\mu,L,\Delta,\sigma$ for all $t$), the same inequality applies when the method is started at any time $s\geq 0$ with initial point $\theta_{s}$ and initial minimizer $\bm{\theta}_{s}^{\star}$, provided the momentum buffer is reset at time $s$. Concretely, if we set $v_{s}=0$ (equivalently $\bm{\theta}_{s-1}=\bm{\theta}_{s}$), and run for $T$ steps with constant stepsize $\gamma$, then (96) (with $t$ replaced by T1T-1 and $(\bm{\theta}_{0},\bm{\theta}_{0}^{\star})$ replaced by $(\bm{\theta}_{s},\bm{\theta}_{s}^{\star})$) yields:
$\mathbb{E}\|\bm{\theta}_{s+T}-\bm{\theta}_{s+T}^{\star}\|^{2}\;\leq\;C_{\beta}\exp\!\Big(-\frac{\mu\gamma}{1-\beta}T\Big)\,\mathbb{E}\|\bm{\theta}_{s}-\bm{\theta}_{s}^{\star}\|^{2}\;+\;\mathcal{E}_{\beta}(\gamma).$ (97)
Proof of (i). Apply (96) with $\gamma=\gamma_{\beta}^{\star}$:
If
$X_{k}:=\bm{\theta}_{t_{k}},\qquad X_{k}^{\star}:=\bm{\theta}_{t_{k}}^{\star}.$
By construction, at the start of each epoch $k$ we restart the momentum buffer (set $v_{t_{k}}=0$), so we may apply the one-epoch bound (97) with $s=t_{k}$, $T=T_{k}$, and $\gamma=\gamma_{k}$:
$\mathbb{E}\|X_{k+1}-X_{k+1}^{\star}\|^{2}\;\leq\;C_{\beta}\exp\!\Big(-\frac{\mu\gamma_{k}}{1-\beta}T_{k}\Big)\,\mathbb{E}\|X_{k}-X_{k}^{\star}\|^{2}\;+\;\mathcal{E}_{\beta}(\gamma_{k}).$ (98)
$T_{k}\;\geq\;\frac{1-\beta}{\mu\gamma_{k}}\log(4C_{\beta}),$
and therefore
Substituting this into (98) yields, for all $k\geq 1$,
$\mathbb{E}\|X_{k+1}-X_{k+1}^{\star}\|^{2}\;\leq\;\frac{1}{4}\,\mathbb{E}\|X_{k}-X_{k}^{\star}\|^{2}\;+\;\mathcal{E}_{\beta}(\gamma_{k}).$ (99)
Base epoch ($k=0$). By the definition of $T_{0}$,
so
Plugging into (98) at $k=0$ gives
$\mathbb{E}\|X_{1}-X_{1}^{\star}\|^{2}\;\leq\;\frac{1}{2}\,\mathcal{E}_{\beta}(\gamma_{0})\;+\;\mathcal{E}_{\beta}(\gamma_{0})\;=\;\frac{3}{2}\,\mathcal{E}_{\beta}(\gamma_{0})\;\leq\;2\,\mathcal{E}_{\beta}(\gamma_{0}).$ (100)
Induction. We claim that for all $k\geq 1$,
$\mathbb{E}\|X_{k}-X_{k}^{\star}\|^{2}\;\leq\;2\,\mathcal{E}_{\beta}(\gamma_{k-1}).$ (101)
The base case $k=1$ follows from (100). Assume (101) holds for some $k\geq 1$. Applying (99) yields
so in particular
Therefore,
which proves (101) at $k+1$ and closes the induction. Hence $\mathbb{E}\|X_{K}-X_{K}^{\star}\|^{2}\leq 2\mathcal{E}_{\beta}(\gamma_{K-1})$. By the definition of $K$,
Thus,
and therefore
Therefore,
@@
Appendix D Proofs of tracking error bounds with high probability
For this section, we will assume a light-tailed assumption on the gradient noise (˜D.1)
Assumption D.1 (Conditional sub-Gaussian gradient noise along iterates).
There exists a constant $\sigma>0$ such that for all $t\geq 0$, $\big\|\bm{\xi}_{t+1}(\bm{\theta}_{t})\mid\mathcal{F}_{t}\big\|_{\Psi_{2}}\leq\sigma$ and $\big\|\bm{\xi}_{t+1}(\bm{\psi}_{t})\mid\mathcal{F}_{t}\big\|_{\Psi_{2}}\leq\sigma\;\;\text{a.s.}$
D.1 Proof for SGD high probability tracking error bound
@@
We will prove the following high probability bound holds for the tracking error in (SGD):
Theorem D.1 (High probability tracking error bound for (SGD)).